Bonds of Zello Corporation with a par value of $1,000 sell for $960, mature in five years, and have a 7% annual coupon rate paid semiannually. Compute the current yield and YTM. Hint: In what situation a bond can have current yield greater than YTM?
Answer #1

Current yield =Annual coupon/current price =70/960 =7.29%

YTM:

=RATE(nper,pmt,pv,fv)

=RATE(5*2,7%/2*1000,-960,1000)*2

=7.99%